INAF LBC+MODS queue observing and AO PISCES, 2012 June 19-20

Observer: C. Arcidiacono, R. Briguglio, K. Boutsia, M. Faccini, F. Vitali
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LBC + PISCES

Summary

Clear skies at the beginning of the night and the average wind is 10m/s.
At 6:00UT the seeing on the dimm is still 1.3"-1.4", so we give up on pursuing Nonino with LBC.
The seeing was variable, with possible thin cirrus passing through, so closed-loops were not as stable as the previous nights.
We close the dome at 10:33UT because of ashes in the air.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Fiore Mrk231 Ks, J ~10min x filter + sky we are up to 75% of requested time
Mesa Hip85157 H, FeII 1000frames in 1hr this is an ADI project - COMPLETED
Testi V1318 Ks, Br-g, H2 10min for Ks and ~20min for Brg/H2 COMPLETED

Details

02:00 H2 flats with Pisces
  • ATTENTION: after image pisces.20120620.0079.fits the light in the dome were ON!

03:38 GCS seeing estimate ~1.5", so we go on with Pisces for now.
  • dimm seeing ~2"

03:51 first preset sent after starting the AOS
  • generic system error ... (?)

04:06 We go to zenith and need to power cycle the W in AGW2

04:22 CCD47 is now ready and we send a preset
  • dimm seeing 1.5" - GCS seeing 1.3"

04:32 shell ripped!

04:35 recovered - send preset again

04:42 start science block for Mrk231

05:02 dimm seeing 1.2" - GCS seeing 1.0"

05:22 during the offset for the sky frames the loop went in pause but the script exited and we were not to recover the loop once the script tried to offset back to the source
  • manually open the loop and resend the preset
  • dimm seeing 1.4" - GCS seeing 1.1"

05:37 The loop is open - resend the preset
  • the loop is open again during the first offset so we move on to the next target

05:46 Slewing to Hip85157

05:52 start collecting ADI data

06:19 loop open - resend the preset
  • the first time the loop gave "safe skip" and runAO failed
  • send preset again

06:40 Since the loo did not seem very stable we optimized the gain, between images pisces*.0784-0793.fits

07:13 loop open - resend preset

07:31 the loop is open - end of ADI
  • send a preset not for ADI but for a not saturated FeII image.

07:34 send preset again for FeII
  • the loop is open again...

07:42 resend preset for not saturated FeII images

07:49 the loop is open again

07:44 dimm seeing 2.1" - problems closing the loop

08:07 still trying to manually close the loop...

08:12 success - loop closed with 100 modes

08:16 it did not last - open loop again
  • loop closed - obtaining calibration data for the photometry

08:33 slew to Testi

08:45 start obtaining science data (V1318)

08:48 loop open - close manually - go on with the script

08:59 loop open during offset for sky ...

09:01 shell rip!!

09:03 shell set - send preset
  • preset failed - the TO has to correct the pointing

09:06 send preset again
  • 09:12 - loop closed
  • start obtaining data again

09:26 loop opened because of very variable seeing - possible cirrus
  • send preset again - problems closing the loop

09:39 loop is closed - start obtaining data - the "not saturated" frames

10:06 trying to manually optimize the gain
  • go on with the science block

10:35 we need to close because there are ashes in the air - End of Night
  • start dark script

-- KonstantinaBoutsia - 20 Jun 2012
Topic revision: r4 - 22 Jun 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback